Suppose favorable weather increased the supply of tomatoes so that the price of tomatoes fell from US$45 a ton to US$35 a ton and quantity sold increased from 300 tons to 420 tons. What is the value of the price elasticity of demand? Show your work.
1
Expert's answer
2020-11-23T05:41:55-0500

Initial price "P_1= 45"

Final price "P_2= 35"

Initial quantity "Q_1= 300\;tons"

Final quantity "Q_2= 420\;tons"

Rise in quantity sold indicates rise in quantity demanded.

Arc elasticity is given as the avg percentage change in quantity demanded/avg percentage change in price.

Midpoint of "Q (Q')= \frac{Q_1+Q_2}{2} = \frac{720}{2} = 360"

Midpoint of "P (P') = \frac{P_1+ P_2}{2} = \frac{80}{2} = 40"

Avg percentage change in quantity "= \frac{Q_2- Q_1}{Q'} = \frac{120}{360} = 0.333"

Avg percentage change in price "= \frac{P_2 - P_1}{P'} = \frac{-10}{40} = -0.25"

Elasticity of demand "= \frac{0.333}{-0.25} = -1.332"

The negative sign indicates inverse relation of price and quantity.